OpenStudy (cicilybailey14):

Help! Find the slope of the line through the given points. (6, -6), (-3, -12) -5/4 -1/2 2/3 -4/5

OpenStudy (anonymous):

graph it

OpenStudy (anonymous):

plot the points

OpenStudy (igreen):

Use the slope formula.. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) Plug in the points: \(m = \dfrac{-12+6}{-3-6}\) Simplify that @Cicilybailey14

OpenStudy (igreen):

Can you add -12 + 6? @Cicilybailey14

OpenStudy (anonymous):

it must be 21

OpenStudy (marvin456):

yea 21